# ANTS: Shaping the Adaptive Negative Textual Space by MLLM for OOD Detection

### v1

@@ Abstract
The introduction of negative labels (NLs) has proven effective in enhancing Out-of-Distribution (OOD) detection. However, existing methods often lack an understanding of OOD images, making it difficult to construct an accurate negative space. In addition, the presence of false negative labels significantly degrades their near-OOD performance. To address these issues, we propose shaping an Adaptive Negative Textual Space (ANTS) by leveraging the understanding and reasoning capabilities of multimodal large language models (MLLMs). Specifically, we identify images likely to be OOD samples as negative images and prompt the MLLM to describe these images, generating expressive negative sentences that precisely characterize the OOD distribution and enhance far-OOD detection. For the near-OOD setting, where OOD samples resemble the in-distribution (ID) subset, we first identify the subset of ID classes that are visually similar to negative images and then leverage the reasoning capability of MLLMs to generate visually similar negative labels tailored to this subset, effectively reducing false negatives and improving near-OOD detection. To balance these two types of negative textual spaces, we design an adaptive weighted score that enables the method to handle different OOD task settings (near-OOD and far-OOD) without relying on task-specific prior knowledge, making it highly adaptable in open environments. On the ImageNet benchmark, our ANTS significantly reduces the FPR95 by 4.2%, establishing a new state-of-the-art. Furthermore, our method is training-free and zero-shot, enabling high scalability.
@@
Figure 1: T-SNE visualization of the ID and OOD image features, the text features of NegLabel [31], EOE [6], OOD ground-truth, and the expressive negative sentences (ENS) of ANTS. We select ImageNet and SUN as the ID and OOD datasets, respectively. NegLabel and EOE lack a good understanding of OOD images, resulting in a greater distance between the OOD images and the text features. In contrast, our ANTS utilizes the MLLMs to understand OOD images during ENS generation, reducing the distance between ENS and OOD images and improving OOD detection performance.
@@
Figure 3: The overall framework of our ANTS. We introduce two negative textual spaces of expressive negative sentences and visually similar negative labels by prompting the MLLM with customized prompts and mined negative images. These two textual spaces are combined with an adaptive weighted score to flexibly handle diverse task settings in open environments.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable performance in classifying test samples that fall into the training distribution [34, 21, 16]. However, it is well-known that DNNs tend to misclassify test samples from unknown classes, which are often called out-of-distribution (OOD) data [22]. Unfortunately, OOD data are inevitably encountered in open environments. Therefore, how to effectively identify OOD data is crucial for the reliable deployment of DNN models in open-world scenarios.
Traditional OOD detection methods in the image domain primarily rely on visual modality information [25, 73, 27, 63, 70]. For example, MSP [22] utilizes the maximum softmax probability of a pre-trained vision model to detect OOD images. Recently, multimodal knowledge has attracted increasing attention in OOD detection [48, 51, 42, 53, 4, 32, 85, 78, 84]. In particular, NegLabel [32] introduces negative labels (NLs) by mining words that are semantically distant from in-distribution (ID) labels, and identifies OOD images by comparing their similarities to NLs and ID labels. Similar approaches generate NLs by prompting LLMs [6] or modifying superclass names [8]. Although these methods have achieved promising performance, they suffer from three key limitations. First, due to the lack of understanding of OOD images, the NLs are positioned far from the OOD image, as illustrated in Fig. 1. Second, these methods struggle with the challenging near-OOD setting, where OOD samples are semantically close to ID labels. NegLabel focuses on generating NLs that are semantically distant from ID classes, inherently overlooking such cases. While EOE [6] introduces visually similar labels for all ID classes to address this problem, it neglects the fact that OOD samples are typically similar to only a subset of ID classes, resulting in many false negative labels (see Fig. 6(a)). Third, these methods rely on the strong assumption that the target task setting (e.g., near OOD or far OOD) is known in advance, allowing for the tailored design of NL generation rules. However, this assumption limits their applicability in complex, unknown, and dynamically changing open environments.
@@ 1 Introduction
To address these challenges, we propose to shape an Adaptive Negative Textual Space (ANTS) by harnessing the understanding and reasoning capabilities of multimodal large language models (MLLMs) [45, 39, 2, 3], as shown in Fig. 2. Specifically, we introduce expressive negative sentences (ENS), which effectively capture fine-grained details of OOD images. These negative sentences are generated by prompting MLLMs to describe online-mined negative images, leveraging their multimodal understanding capabilities and significantly enhancing the traditional far-OOD detection. While ENS shows greater expressive power in identifying far-OOD samples, it faces challenges in handling the near-OOD setting, where OOD samples are semantically close to certain ID classes. To address this, we dynamically identify the subset of ID classes most similar to the negative images and utilize the reasoning capabilities of MLLMs to construct visually similar negative labels (VSNL) tailored for this subset. This targeted approach reduces false negative labels and improves performance in the near-OOD setting (see Fig. 6(a)). Finally, to ensure adaptability across diverse task settings in open environments, we introduce an adaptive weighted score function to balance the two distinct negative textual spaces. This dynamic mechanism enables the model to seamlessly handle both near-OOD and far-OOD scenarios without prior knowledge of the task settings. The overall framework is presented in Fig. 3.
@@
We conduct extensive experiments to validate the advantages of our ANTS method. On the large-scale ImageNet dataset, our approach significantly reduces FPR95 by 4.2% and 3.25% in the far-OOD and challenging near-OOD detection settings, respectively. Moreover, our method operates in a zero-shot and training-free manner, demonstrating strong scalability across different MLLMs. We summarize our contributions as follows:
@@
  • To overcome these limitations, we propose the ANTS approach by leveraging the understanding and reasoning capabilities of MLLMs. Specifically, we (1) design two types of prompt for MLLMs to generate expressive negative sentences and visually similar negative labels; (2) introduce two strategies including Negative Images Mining and Visually Similar ID-Classes Mining to avoid interference from ID noise and generate false negative labels; and (3) design an adaptive weighted score to dynamically balance these two text spaces in open environments.
  • Extensive experiments are conducted to validate the proposed components. Our method demonstrates new state-of-the-art performance on both near-OOD and far-OOD detection tasks, achieving significant gains without any prior knowledge about the task setting.
@@ 2 Related Work
Traditional OOD Detection. Traditional OOD detection methods can be categorized into the following groups: (1) classification-based methods [22, 43, 36, 46, 60, 83, 63, 15, 62, 44, 55, 31, 66, 80, 74, 26, 23, 19, 79, 54, 49] that distinguish ID and OOD samples by designing a score function; (2) density-based methods [89, 58, 1, 56, 28, 71] that detect OOD samples by evaluating the likelihood or density of test data derived from probabilistic models; (3) distance-based methods [81, 67, 50] that classify test samples as OOD if they deviate significantly from the prototypes of in-distribution classes; (4) reconstruction-based methods [13, 87, 76, 29, 38] that focus on detecting OOD samples by analyzing the reconstruction errors within a reconstruction framework. However, traditional methods primarily rely on visual information and often overlook the integration of textual information, limiting their ability to fully leverage multimodal data for OOD detection.
OOD Detection with Vision Language Model. VLM-based OOD detection methods can be broadly categorized into two settings: few-shot and zero-shot. Few-shot methods enhance OOD detection by using negative prompts to define boundaries between ID and OOD images  [53, 42, 4], or by integrating non-ID or local ID regions for regularization  [51, 35]. For zero-shot OOD detection, some works [48, 84, 77, 33, 88] design post-hoc strategies that utilize softmax scores or image feature information during testing. Some methods [72, 85] leverage auxiliary datasets to strengthen the detection of OOD samples. Other approaches [18, 55, 32, 6, 8] retrieve negative labels from corpus databases or generating them using LLMs. However, these NLs methods lack understanding of actual OOD images, the semantic gap with OOD images limits their OOD detection capabilities.
Multimodal Large Language Models. MLLMs [45, 39, 2, 3, 41, 37] demonstrate strong capabilities across various visual tasks. By integrating large language models, these systems can better interpret instructions and more accurately complete the corresponding tasks. Some approaches [30, 11, 7, 9] have leveraged the understanding and reasoning capabilities of MLLMs for industrial anomaly detection, utilizing their ability to capture fine-grained anomaly semantics. While MLLMs have proven effective in industrial anomaly detection, no prior work has investigated their application to large-scale OOD detection. Our approach is the first to exploit MLLMs’ test time understanding and reasoning capabilities to construct an adaptive and expressive negative textual space, significantly enhancing OOD detection performance.
@@ 3 Preliminary
OOD Detection Setup. Denote by $\mathcal{X}$ the image space and $\mathcal{Y}=\{y_{1},\dots,y_{N}\}$ the ID label space, with examples $\mathcal{Y}=\{cat,dog,\dots,bird\}$ and $N$ denoting the total number of classes. Given ${\bm{x}}_{in}\in\mathcal{X}$ as the ID random variable and ${\bm{x}}_{ood}\in\mathcal{X}$ as the OOD random variable, we denote their respective distributions as $\mathcal{P}_{{\bm{x}}_{in}}$ and $\mathcal{P}_{{\bm{x}}_{ood}}$. In closed-set scenarios, a test image $\bm{x}$} is expected to belong to one ID class, i.e., ${\bm{x}}\in\mathcal{P}_{{\bm{x}}_{in}}$ and $y\in\mathcal{Y}$, where $y$ is the label of $\bm{x}$}. However, in real-world scenarios, AI systems may encounter samples that do not match any known class, i.e., ${\bm{x}}\in\mathcal{P}_{{\bm{x}}_{ood}}$ and $y\notin\mathcal{Y}$, resulting in potential misclassifications and safety concerns [61, 52]. To tackle these issues, OOD detection aims to distinguish ID and OOD samples using a scoring function $S$:
@@
OOD Detection with NLs. Enhancing OOD detection with textual knowledge has recently garnered increasing attention [48, 72, 84], while a representative type of approach introduces NLs [32, 6]. Specifically, in addition to the ID labels $\mathcal{Y}$, these methods introduce a disjoint set of NLs $\mathcal{Y}^{-}$ and classify a test sample as OOD if it exhibits high similarity to NLs and low similarity to ID labels. In this process, the quality of NLs is crucial. The pioneering method, NegLabel [32], selects words with large cosine distance to ID labels in a large corpus dataset $\mathcal{Y}^{c}=\left\{\widetilde{y}_{1},\widetilde{y}_{2},\ldots,\widetilde{y}_{K}\right\}$ as NLs:
@@
where the CLIP-like model $f_{clip}$ defines the label similarity space. $K$ and $M$ represent the numbers of candidate labels in $\mathcal{Y}^{c}$ and the selected NLs in $\mathcal{Y}^{-}_{nl}$, where $M\leq K$. Another representative work, EOE [6], uses prompts to guide an LLM to generate NLs:
@@
where $\rho_{neg}$ is a carefully designed textual prompt for the LLM $f_{llm}$. Given the generated NLs (e.g., $\mathcal{Y}^{-}_{nl}$ [32]), the score function for OOD detection can be formulated as:
@@ 4.1 Motivation
Although NegLabel [32] and EOE [6] have advanced OOD detection using NLs, they face three key limitations: (1) lacking of understanding of OOD images; (2) poor performance in near-OOD settings due to false negatives by neglecting visually similar classes; and (3) reliance on prior task knowledge, limiting adaptability in open environments. This motivates us to raise the following question: Can we leverage the test time understanding and reasoning capabilities of MLLMs to shape a more accurate and comprehensive negative textual space?
In this work, we attempt to answer this question by designing different prompts for MLLMs to shape a more accurate negative textual space, as shown in Fig. 2. The overall pipeline of our method is illustrated in Fig. 3.
@@ 4.2 Expressive Negative Sentences
Negative Images Mining. We leverage the image understanding capabilities of MLLMs to generate expressive negative sentences by describing negative images, which are test images likely to be OOD samples. We identify these negative images using the OOD detector of NegLabel, where test images with $\mathcal{S}_{nl}({\bm{x}})<\gamma$ are selected as negative images:
$\mathcal{X}_{neg}=\{{\bm{x}}\mid S_{nl}({\bm{x}})<\gamma,\ {\bm{x}}\in\mathcal{X}_{test}\},$ (5)
where $\mathcal{X}_{test}$ denotes the test data. We find that manually defining a fixed $\gamma$ is challenging for handling different testing scenarios, as the optimal threshold varies between near and far OOD settings, as analyzed in Fig. 6(b). To address this issue, we develop an adaptive threshold determination strategy based on the characteristics of the test data. Specifically, we first filter out test samples with high $\mathcal{S}_{nl}$ scores using Eq. 5, as these samples are highly likely to be ID samples. For the remaining negative images $\widehat{\mathcal{X}}_{neg}$, which fall into a mixed set of ID and OOD samples, we select a proportion $\eta$ of images with the lowest $S_{nl}$ scores, and the adaptive threshold $\gamma^{*}$ can be formulated as:
@@
(b) Different thresholds $\gamma$
(c) $S_{ens}({\bm{v}})$ Vs. $S_{vsnl}({\bm{v}})$
Figure 6: (a) Our VSNL generates visually similar labels only for the ID class subset, whose images are most similar to the near OOD samples, largely reducing false negative labels. (b) Different task settings prefer different thresholds, and our proposed method can adaptively mine negative images, implicitly setting an adaptive threshold. (c) $S_{ens}$ and $S_{vsnl}$ perform differently on far and near OOD, providing clues for designing an adaptive weight $\lambda$ in Eq. 14.

### v3

# Wenjie Zhu1,2  Yabin Zhang3footnotemark:  Xin Jin2  Wenjun Zeng2  Lei Zhang1footnotemark:

@@ Abstract
The introduction of negative labels (NLs) has proven effective in enhancing Out-of-Distribution (OOD) detection. However, existing methods often lack an understanding of OOD images, making it difficult to construct an accurate negative space. Furthermore, the absence of negative labels semantically similar to ID labels constrains their capability in near-OOD detection. To address these issues, we propose shaping an Adaptive Negative Textual Space (ANTS) by leveraging the understanding and reasoning capabilities of multimodal large language models (MLLMs). Specifically, we cache images likely to be OOD samples from the historical test images and prompt the MLLM to describe these images, generating expressive negative sentences that precisely characterize the OOD distribution and enhance far-OOD detection. For the near-OOD setting, where OOD samples resemble the in-distribution (ID) subset, we cache the subset of ID classes that are visually similar to historical test images and then leverage MLLM reasoning to generate visually similar negative labels tailored to this subset, effectively reducing false negatives and improving near-OOD detection. To balance these two types of negative textual spaces, we design an adaptive weighted score that enables the method to handle different OOD task settings (near-OOD and far-OOD), making it highly adaptable in open environments. On the ImageNet benchmark, our ANTS significantly reduces the FPR95 by 3.1%, establishing a new state-of-the-art. Furthermore, our method is training-free and zero-shot, enabling high scalability.
@@
Figure 1: T-SNE visualization of the ID and OOD image features, the text features of NegLabel [jiang2023detecting], EOE [cao2024envisioning], OOD ground-truth, and the expressive negative sentences (ENS) of ANTS. We select ImageNet and SUN as the ID and OOD datasets, respectively. NegLabel and EOE lack a good understanding of OOD images, resulting in a greater distance between the OOD images and the text features. In contrast, our ANTS utilizes the MLLMs to understand OOD images during ENS generation, reducing the distance between ENS and OOD images and improving OOD detection performance.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable performance in classifying test samples that fall into the training distribution [krizhevsky2009learning, he2016deep, dosovitskiy2020image]. However, it is well-known that DNNs tend to misclassify test samples from unknown classes, which are often called out-of-distribution (OOD) data [hendrycks2016baseline]. Unfortunately, OOD data are inevitably encountered in open environments. Therefore, how to effectively identify OOD data is crucial for the reliable deployment of DNN models in open-world scenarios.
Traditional OOD detection methods in the image domain primarily rely on visual modality information [hsu2020generalized, wang2021energy, huang2021importance, sun2021react, wang2021can]. For example, MSP [hendrycks2016baseline] utilizes the maximum softmax probability of a pre-trained vision model to detect OOD images. Recently, multimodal knowledge has attracted increasing attention in OOD detection [ming2022delving, miyai2024locoop, li2024learning, nie2024out, bai2024id, jiang2024negative, zhang2024lapt, yu2024self, zhang2024adaneg]. In particular, NegLabel [jiang2024negative] introduces negative labels (NLs) by mining words that are semantically distant from in-distribution (ID) labels, and identifies OOD images by comparing their similarities to NLs and ID labels. Similar approaches generate NLs by prompting LLMs [cao2024envisioning] or modifying superclass names [chen2024conjugated]. Although these methods have achieved promising performance, they suffer from three key limitations. First, due to the lack of understanding of OOD images, the NLs are positioned far from the OOD image, as illustrated in Fig. 1. Second, these methods struggle with the challenging near-OOD setting, where OOD samples are semantically close to ID labels. NegLabel focuses on generating NLs that are semantically distant from ID classes, inherently overlooking such cases. While EOE [cao2024envisioning] introduces visually similar labels for all ID classes to address this problem, it neglects the fact that OOD samples are typically similar to only a subset of ID classes, resulting in many false negative labels (see Fig. 6(a)). Third, these methods rely on the strong assumption that the target task setting (e.g., near OOD or far OOD) is known in advance, allowing for the tailored design of NL generation rules. However, this assumption limits their applicability in complex, unknown, and dynamically changing open environments.
@@ 1 Introduction
Figure 3: The overall framework of our ANTS. ANTS framework consists of in three stages: (1) caching negative images and visually similar ID classes mined from historical test images; (2) shaping two negative textual spaces by prompting an MLLM with the cached data to generate expressive negative sentences and visually similar labels; and (3) performing online evaluation of the test image using an adaptively weighted combination of these textual spaces.
@@
To address these challenges, we propose to shape an Adaptive Negative Textual Space (ANTS) by harnessing the understanding and reasoning capabilities of multimodal large language models (MLLMs) [liu2023visual, li2023blip, achiam2023gpt, bai2025qwen2], as shown in Fig. 2. Specifically, we introduce expressive negative sentences (ENS), which effectively capture fine-grained details of OOD images. These negative sentences are generated by prompting MLLMs to describe online-mined negative images, leveraging their multimodal understanding capabilities and significantly enhancing the traditional far-OOD detection. While ENS shows greater expressive power in identifying far-OOD samples, it faces challenges in handling the near-OOD setting, where OOD samples are semantically close to certain ID classes. To address this, we dynamically identify the subset of ID classes most similar to the negative images and utilize the reasoning capabilities of MLLMs to construct visually similar negative labels (VSNL) tailored for this subset. This targeted approach reduces false negative labels and improves performance in the near-OOD setting (see Fig. 6(a)). Finally, to ensure adaptability across diverse task settings in open environments, we introduce an adaptive weighted score function to balance the two distinct negative textual spaces. This dynamic mechanism enables the model to seamlessly handle both near-OOD and far-OOD scenarios without prior knowledge of the task settings. The overall framework is presented in Fig. 3.
@@
We conduct extensive experiments to validate the advantages of our ANTS method. On the large-scale ImageNet dataset, our approach significantly reduces FPR95 by 3.1% and 3.25% in the far-OOD and challenging near-OOD detection settings, respectively. Moreover, our method operates in a zero-shot and training-free manner, demonstrating strong scalability across different MLLMs. We summarize our contributions as follows:
@@
  • To overcome these limitations, we propose the ANTS approach by leveraging the understanding and reasoning capabilities of MLLMs. Specifically, we (1) introduce two strategies including Negative Images Mining and Visually Similar ID-Classes Mining to avoid interference from ID noise and generate false negative labels; (2) design two types of prompt for MLLMs to generate expressive negative sentences and visually similar negative labels; and (3) design an adaptive weighted score to dynamically balance these two text spaces in open environments.
  • Extensive experiments are conducted to validate the proposed components. Our method demonstrates new state-of-the-art performance on both near-OOD and far-OOD detection tasks. Our method is training-free, zero-shot, and does not require any auxiliary outlier images.
@@ 2 Related Work
Traditional OOD Detection. Traditional OOD detection methods can be categorized into the following groups: (1) classification-based methods [hendrycks2016baseline, liang2017enhancing, lee2018simple, liu2020energy, sastry2020detecting, zhang2022out, sun2021react, dong2022neural, sun2022dice, lin2021mood, park2023nearest, jiang2023detecting, thulasidasan2019mixup, yun2019cutmix, wei2022mitigating, huang2021mos, hendrycks2018deep, fort2021exploring, yu2019unsupervised, papadopoulos2021outlier, ming2022poem] that distinguish ID and OOD samples by designing a score function; (2) density-based methods [zong2018deep, ren2019likelihood, abati2019latent, pidhorskyi2018generative, jiang2021revisiting, wang2022vim] that detect OOD samples by evaluating the likelihood or density of test data derived from probabilistic models; (3) distance-based methods [zaeemzadeh2021out, van2020uncertainty, ming2022exploit] that detect OOD samples by measuring their deviation from in-distribution class prototypes.
OOD Detection with Vision Language Model. VLM-based OOD detection methods can be broadly categorized into two settings: few-shot and zero-shot. Few-shot methods enhance OOD detection by using negative prompts to define boundaries between ID and OOD images [nie2024out, li2024learning, bai2024id], or by integrating non-ID or local ID regions for regularization [miyai2024locoop, lafon2024gallop]. For zero-shot OOD detection, some works [ming2022delving, zhang2024adaneg, yang2025oodd, kim2025enhanced, zhu2025knowledge] design post-hoc strategies that utilize softmax scores or image feature information during testing. Some methods [wang2023clipn, zhang2024lapt] leverage auxiliary datasets to strengthen the detection of OOD samples. Other approaches [esmaeilpour2022zero, park2023nearest, jiang2024negative, cao2024envisioning, chen2024conjugated, dai2023exploring] retrieve negative labels from corpus databases or generating them using LLMs. However, these NLs methods lack understanding of actual OOD images, the semantic gap with OOD images limits their OOD detection capabilities.
@@ 3 Preliminary
OOD Detection Setup. Denote by $\mathcal{X}$ the image space and $\mathcal{Y}=\{y_{1},\dots,y_{N}\}$ the ID label space, with examples $\mathcal{Y}=\{cat,dog,\dots,bird\}$ and $N$ denoting the total number of classes. Given ${\bm{x}}_{in}\in\mathcal{X}$ as the ID random variable and ${\bm{x}}_{ood}\in\mathcal{X}$ as the OOD random variable, we denote their respective distributions as $\mathcal{P}_{{\bm{x}}_{in}}$ and $\mathcal{P}_{{\bm{x}}_{ood}}$. In closed-set scenarios, a test image $\bm{x}$} is expected to belong to one ID class, i.e., ${\bm{x}}\in\mathcal{P}_{{\bm{x}}_{in}}$ and $y\in\mathcal{Y}$, where $y$ is the label of $\bm{x}$}. However, in real-world scenarios, AI systems may encounter samples that do not match any known class, i.e., ${\bm{x}}\in\mathcal{P}_{{\bm{x}}_{ood}}$ and $y\notin\mathcal{Y}$, resulting in potential misclassifications and safety concerns [scheirer2012toward, nguyen2015deep]. To tackle these issues, OOD detection aims to distinguish ID and OOD samples using a scoring function $S$:
@@
OOD Detection with NLs. Enhancing OOD detection with textual knowledge has recently garnered increasing attention [ming2022delving, wang2023clipn, zhang2024adaneg], while a representative type of approach introduces NLs [jiang2024negative, cao2024envisioning]. Specifically, in addition to the ID labels $\mathcal{Y}$, these methods introduce a disjoint set of NLs $\mathcal{Y}^{-}$ and classify a test sample as OOD if it exhibits high similarity to NLs and low similarity to ID labels. In this process, the quality of NLs is crucial. The pioneering method, NegLabel [jiang2024negative], selects words with large cosine distance to ID labels in a large corpus dataset $\mathcal{Y}^{c}=\left\{\widetilde{y}_{1},\widetilde{y}_{2},\ldots,\widetilde{y}_{K}\right\}$ as NLs:
@@
where the CLIP-like model $f_{clip}$ defines the label similarity space. $K$ and $M$ represent the numbers of candidate labels in $\mathcal{Y}^{c}$ and the selected NLs in $\mathcal{Y}^{-}_{nl}$, where $M\leq K$. Another representative work, EOE [cao2024envisioning], uses prompts to guide an LLM to generate NLs:
@@
where $\rho_{neg}$ is a carefully designed textual prompt for the LLM $f_{llm}$. Given the generated NLs (e.g., $\mathcal{Y}^{-}_{nl}$ [jiang2024negative]), the score function for OOD detection can be formulated as:
@@ 4.1 Motivation
Although NegLabel [jiang2024negative] and EOE [cao2024envisioning] have advanced OOD detection using NLs, they face three key limitations: (1) lacking of understanding of OOD images, as shown in Fig.1; (2) poor performance in near-OOD settings due to false negatives by neglecting visually similar classes; and (3) reliance on prior task knowledge, limiting adaptability in open environments. This motivates us to raise the following question:
Can we leverage the test time understanding and reasoning capabilities of MLLMs to shape a more accurate and comprehensive negative textual space?
In this work, we attempt to answer this question by designing different prompts for MLLMs to leverage their test-time understanding and reasoning capabilities for OOD detection, as shown in Fig.2. The overall pipeline of our method is illustrated in Fig. 3.
@@ 4.2 Expressive Negative Sentences
Negative Images Mining. We leverage the image understanding capabilities of MLLMs to generate expressive negative sentences by describing negative images, which are historical test images likely to be OOD samples. We identify these negative images using the OOD detector of NegLabel, where historical test images with $\mathcal{S}_{nl}({\bm{x}})<\gamma$ are selected as negative images:
$\mathcal{X}_{neg}=\{{\bm{x}}\mid S_{nl}({\bm{x}})<\gamma,\ {\bm{x}}\in\mathcal{X}^{\rm his}_{test}\},$ (5)
where $\mathcal{X}^{\rm his}_{test}$ denotes the historical test data. We find that manually defining a fixed $\gamma$ is challenging for handling different testing scenarios, as the optimal threshold varies between different OOD datasets, as analyzed in Fig. 6(b). To address this issue, we develop an adaptive threshold determination strategy based on the characteristics of the historical test data. Specifically, we filter out historical test samples with high $\mathcal{S}_{nl}$ scores using Eq. 5, as these samples are highly likely to be ID samples. For the remaining negative images $\widehat{\mathcal{X}}_{neg}$, which fall into a mixed set of ID and OOD samples, we select a proportion $\eta$ of images with the lowest $S_{nl}$ scores, and the adaptive threshold $\gamma^{*}$ can be formulated as:
@@
(b) Different optimal thresholds $\gamma$
(c) $1-S_{ens}({\bm{v}})$ Vs. $1-S_{vsnl}({\bm{v}})$
Figure 6: (a) Our VSNL generates visually similar labels only for the ID class subset, whose images are most similar to the near OOD samples, largely reducing false negative labels. (b) Different OOD datasets prefer different thresholds, and our proposed method can cache the historical test images and adaptively mine negative images, implicitly setting an dataset adaptive threshold. (c) $S_{ens}$ and $S_{vsnl}$ perform differently on far and near OOD, providing clues for designing an adaptive weight $\lambda$ in Eq. 14.